OpenStudy (anonymous):

What is the slope for the following table of values? The table represents the total amount spent in a paperback book club after an introductory fee of $12. # books purchased 1, 2, 3, 4 Cost of book club 21, 30, 39, 48 A. 3 B. 5 C. 7 D. 9

Directrix (directrix):

X: # books purchased 1, 2, 3, 4 Y: Cost of book club 21, 30, 39, 48 Slope = change in y over change in x (21 - 30)/ (1-2) = -9(-1) = 9 (48 - 39)/ (4-3) = 9/1 = 9 Slope is 9 -->